OpenStudy (anonymous):

What is the slope of the line whose equation is –48 = 2x − 8y?

OpenStudy (anonymous):

@parker.goodbar

OpenStudy (parker.goodbar):

first We need to isolate y –48 = 2x − 8y –48-2x = -8y -8y = -2x - 48 y = (-2x - 48)/(-8) y = (-2x)/(-8) - 48/(-8) y = (1/4)x + 6 The last equation above is in the form y = mx+b where m = 1/4 is the slope

OpenStudy (parker.goodbar):

you still there @smartnoekrjdnkfnenr